3. A stock has had returns of 44.00 percent, 28.00 percent, 39.00 percent, −16.00 percent, 26.00...

3.

A stock has had returns of 44.00 percent, 28.00 percent, 39.00 percent, −16.00 percent, 26.00 percent, and −58.00 percent over the last six years. What are the arithmetic and geometric returns for the stock? (Do not round intermediate calculations. Round the final answers to 2 decimal places.)

Arithmetic average return %
Geometric average return %

4.

A stock has had returns of 17.62 percent, 12.38 percent, 6.42 percent, 27.94 percent, and −13.94 percent over the past five years, respectively.

What was the holding-period return for the stock? (Do not round intermediate calculations. Round the final answer to 2 decimal places.)

Holding-period return             %
Answer #1

Ans 3.

Arithmetic Average Return = Sum of all the Returns / No of times

= (44.00 + 28.00 + 39.00 + -16.00 + 26.00 + -58.00) / 6 = 63 / 6 = 10.50%

Average Geometric Returns = (1*r x 1*r x ... n)^ 1/6 - 1

=(1.44*1.28*1.39*0.84*1.26*0.42)^(1/6)-1 = 0.0219 or 2.19%

Ans 4)

Holding Period Return = (1+r1)* (1+r2)* (1+rn) -1

= (1.1762 * 1.1238* 1.0642 * 1.2794 * .8606) - 1 = 0.5488 or 54.88%

Hope this helps, let me know in case of any confusion. Feel free to share your feedback. Thanks and have a good day.